# Understanding the Issues, Their Causes and Solutions in Microservices Systems: An Empirical Study

### v2

@@ 3.4 Practitioners’ Perspective (RQ4)
Practitioners’ Perspective on Issues: We asked the microservices participants which issue they faced while developing microservices systems. The majority of the respondents mentioned that they face all of the issues while developing microservices systems. The practitioners’ responses presented in Table V show that IC1: Technical Debt (46.67% Very Often, 24.00% Often), IC2: Continuous Integration and Delivery Issue (26.67% Very Often, 42.67% Often), and IC5: Security Issue (18.67% Very Often, 64.00% Often) occur most frequently than other categories of issues. Some practitioners also suggested 6 types of issues (with 9 instances of issues) that were not part of the initial taxonomy of issues in microservices systems. We added microservices practitioners’ suggested types of issues in Figure 4 and the Issue Taxonomy sheet in [19]. One representative quotation about other types of issues is depicted below.
@@
Practitioners’ Perspective on Causes: Our survey participants also confirmed the identified causes that could lead to issues in microservices systems. The practitioners’ responses presented in Table V show that CC1: General Programming Error (41.33% Strongly Agree, 38.67% Agree), CC8: Fragile Code (23.33% Strongly Agree, 46.67% Agree), and CC2: Missing Features and Artifacts (12.67% Strongly Agree, 55.33% Agree) are the top three cause categories. Some practitioners also suggested 5 types of causes (with 11 instances of causes) that were not part of the initial taxonomy of causes in microservices systems. We added microservices practitioners’ suggested types of causes in Table V and the Cause Taxonomy sheet in [19]. One representative quotation about suggested types of causes is depicted below.
@@
Practitioners’ Perspective on Solutions: We also recorded the practitioners’ perspective about the solutions for the issues occurring during microservices system development. The practitioners’ responses presented in Table V show that SC1: Add Artifacts (42.00% Strongly Agree, 39.33% Agree), SC3: Modify Artifact (27.33% Strongly Agree, 37.33% Agree), and SC6: Manage Configuration and Execution (18.00% Strongly Agree, 48.67% Agree) and are the top three solution categories that have been used to address the microservices issues. One practitioner also suggested one type of solution (with one instance of solution) that was not part of the initial taxonomy of solutions in microservices systems. We added microservices practitioners’ suggested types of solutions in Table V and the Solution Taxonomy sheet in [19]. One representative quotation about the suggested solutions is depicted below.
@@
TABLE V: Practitioners’ perspective (in %) on the issue, cause, and solution categories in microservices systems
@@
Statistical significance on the issue, cause, and solution categories in microservices systems: We analyzed the practitioners’ responses across one pair of demographic groups in Table VI. The first column of Table VI lists the categories of issues, causes, and solutions presented to the survey participants. The subsequent columns of Table VI show the Likert Distribution, Mean, p-value, and Effect Size. Analyzing survey data using the Likert Distribution, Mean, p-value, and Effect Size provides insights into response patterns, central tendency, statistical significance, and practical significance, respectively. The practitioners’ responses are grouped into Experience $\leq$ 6 Years vs. Experience > 6 Years group to check the test whether there are statistical differences between the two groups on the same variable or not.
@@
{tcolorbox}
[sharp corners, boxrule=0.1mm,] Key Findings of RQ4: (i) There are no major statistically significant differences in the issue, cause, and solution categories in practitioners’ responses, (ii) practitioners frequently face Technical Debt, CI/CD, and Security issues, (iii) most causes are associated with General Programming Errors, Fragile Code, and Missing Features and Artifacts, and (iv) most solutions are associated with Add Artifacts, Modify Artifact, and Manage Configuration and Execution. The survey participants generally confirmed the categories of issues, causes, and solutions derived from mining developer discussions in open-source microservices systems. However, the survey participants also indicated several other types of issues, causes, and solutions that help improve our taxonomies.
TABLE VI: Statistical significance on the issue, cause, and solution categories in microservices systems
@@ 5.1 Internal Validity
Improper project selection. The first internal validity threat to our study is an improper selection of OSS projects for executing our research plan. Mitigation. We used a multi-step project selection approach to control the possible threat associated with subject system selection (see Phase 1 in Section 2.2 and Figure 2). A step-wise and criteria-driven approach for the project selection has been used that helped us to include the relevant (see Table I) and eliminate irrelevant open-source microservices projects.
@@
Participants selection and background: The survey participants may not have sufficient expertise to answer questions. Mitigation. We searched for microservices practitioners through personal contacts and relevant platforms (e.g., LinkedIn and GitHub). We explicitly made participants’ characteristics (e.g., roles and responsibilities) in the interview and survey preamble, and we selected only those practitioners with sufficient experience designing, developing, and operationalizing microservices systems. For example, the average experience of the interview participants is 5.33 years, and they mainly have responsibilities for designing and developing microservices systems (see Table II and Figure 3).
@@ 5.2 Construct Validity
Inadequate explanation of the constructs: This threat refers to the fact that the study constructs are not sufficiently described. Mitigation. To deal with this threat, we prepared the protocols for mining developer discussions and conducting interviews and surveys, and these protocols were continuously improved during the internal meetings, feedback, and taxonomy refinements. Mainly, the authors had meetings (i) to establish a common understanding of issues, causes, and solutions (see Figure 1), (ii) for extracting the required data to answer the RQs, and (iii) for evaluating interview and survey question format, understandability, and consistency. We also invited two survey-based research experts to check the validity and integrity of the survey questions. Based on their feedback, we included Figure 4, Table III, and Table IV as a part of the interview and survey questions for improving the participants’ understandability of the issues, their causes and solutions in microservices systems.
@@ 5.3 External Validity
Interview data saturation: This threat arises from the possibility that the 15 interviews we conducted to validate and expand our proposed taxonomies may not be sufficient to achieve data saturation in our study. Mitigation. To address this threat, we conducted interviews with practitioners who had varied backgrounds (i.e., responsibilities, skills in programming languages, work domains, experience, countries) in microservices systems (see Table II). Second, we asked open-ended questions, which allowed the interviewees to spontaneously express their views on the issues, causes, and solutions (see the Interview Questionnaire sheet in [19]). Third, we presented the interviewees with pre-existing taxonomies that were derived from analyzing developer discussions in 15 open-source microservices systems. We asked the interviewees to pinpoint any issue, cause, and solution they felt absent. Fourth, while we were analyzing interview data, we started observing recurring themes and patterns in the responses, which was an indication of data saturation. As we did not encounter any new significant issues, causes, or solutions after the 15th interview, we decided to stop conducting further interviews. However, we encourage further research involving more interviews or additional data sources to provide more insights and refinement to our findings.
@@ 6.1.1 Mining (Anti-)Patterns and Tactics to Architect MSA
From an MSA perspective, architectural patterns [34] and tactics [66] represent empirically derived knowledge, best practices, and recurring solutions to address frequently occurring issues during architecture-centric design and development of service-driven systems [10]. To discover patterns and tactics for architecting MSA, a number of studies have focused on analyzing open source repositories – mining version controls, searching change logs, and exploring design documents etc. – to investigate historical data that represents recurring solutions as patterns [10, 67, 68, 69]. The investigation of historical data involves postmortem analysis of development activities (e.g., software refactoring, testing, evolution) [70] as well as the knowledge of architects (e.g., developer discussions, code documentation) [71], which is captured via open source repositories, such as GitHub, Stack Overflow, or customized databases [10, 67, 71]. Specifically, Marquez et al. [10] explored source code artifacts of microservices projects, i.e., configuration files, and framework dependencies, available on GitHub to mine 17 architectural patterns addressing a set of quality attributes. Compared to the design and implementation-specific knowledge for microservices reported in [10, 67], Armin et al. [68] explored the evolution phase from traditional architectures to MSA and identified 15 migration patterns that support the evolution of legacy software towards a modular MSA. Compared to the pattern-based solutions discussed above, architectural tactics represent design decisions that focus on improving one specific quality aspect of MSA, such as service availability and fault avoidance for security critical systems [66].
@@
While patterns promote best practices to develop MSA, anti-patterns represent a class of patterns that have been perceived as best practices and commonly used but are proven to be ineffective and/or counterproductive, such as bad smells in service design [14]. Taibi et al. [70] conducted interviews of microservices developers, i.e., accumulating practitioners’ experiences to create a taxonomy of 20 anti-patterns including organizational (team oriented and technology/tool oriented) and technical (internal and communication) anti-patterns. The anti-pattern taxonomy aims to help microservices practitioners to avoid counterproductive patterns and design decisions. In comparison to the repository mining approaches that investigate code and architectural-centric artifacts available on GitHub such as [66, 10, 69], Bandeira et al. [71] reflected a human-centric view of MSA design by analyzing developer discussions on Stack Overflow. Their study classified a total of 1,043 microservice tagged posts into three categories namely technical (44.87%), conceptual (30%), and non-related (25.13%) discussions to explore the processes, tools, issues, solutions, etc. that developers find most exciting or challenging while implementing MSA.
@@ 6.3 Conclusive Summary
We provide a comparative analysis of the most relevant existing research and our proposed study in Table VII based on (i) focus of investigation: specific aspects to be empirically investigated, (ii) research method: specific approaches employed for investigation, (iii) number of issues: specific number of the identified issues, and (iv) method for evaluation: specific method for evaluation. Additionally, the study reference, year of publication, and contribution of the research study are provided as complementary information for a fine-grained presentation of the results. For example, an empirical study reported in [15] published in 2021 relies on an industrial survey (i.e., applied research method) to investigate common faults and debugging practices in microservices systems (i.e., focus of investigation). The study results (i.e., main contribution) highlight typical faults, state-of-practice on debugging them, and indicate the challenges faced by developers for microservices systems.
TABLE VII: Comparative analysis and summary of existing research vs. this research
@@
[15]
2021
Fault analysis, Debugging
Industry survey
Not explicit mentioned
@@
[16]
2020
Performance issues
MSA application monitoring
Not explicit mentioned
@@
[72]
2019
Evolvability analysis
Structured interviews
16 types of challenges
@@
[75]
2020
Code smells
Static code analysis
11 types code smells
@@
[78]
2023
Security points
Repository mining
@@
[79]
2022
Microservice problems and solutions
Repository mining
14 problems, 47 solutions
@@
[80]
2022
Architectural smells
Interviews, Repository mining
6 types of architectural smells
@@
Our previous study [18]
2022
Issues and Causes in microservices systems
Repository mining
138 types of issues, 109 types of causes
@@
This study
2023
Issues, Causes, and Solutions in microservices systems
Repository mining (2,641 issues), Structured interviews (15 interviews), Practitioners’ survey (150 responses)
402 types of issues, 228 types of causes, 177 types of solutions
@@ Acknowledgments
This work is partially sponsored by the NSFC with Grant No. 62172311. The authors would also like to thank the participants of the interviews and online survey.

### v3

@@ 3.4 Practitioners’ Perspective (RQ4)
Practitioners’ Perspective on Issues: We asked the microservices participants which issue they faced while developing microservices systems. The majority of the respondents mentioned that they face all of the issues while developing microservices systems. The practitioners’ responses presented in Table 5 show that IC1: Technical Debt (46.67% Very Often, 24.00% Often), IC2: Continuous Integration and Delivery Issue (26.67% Very Often, 42.67% Often), and IC5: Security Issue (18.67% Very Often, 64.00% Often) occur most frequently than other categories of issues. Some practitioners also suggested 6 types of issues (with 9 instances of issues) that were not part of the initial taxonomy of issues in microservices systems. We added microservices practitioners’ suggested types of issues in Figure 4 and the Issue Taxonomy sheet in [19]. One representative quotation about other types of issues is depicted below.
@@
Practitioners’ Perspective on Causes: Our survey participants also confirmed the identified causes that could lead to issues in microservices systems. The practitioners’ responses presented in Table 5 show that CC1: General Programming Error (41.33% Strongly Agree, 38.67% Agree), CC8: Fragile Code (23.33% Strongly Agree, 46.67% Agree), and CC2: Missing Features and Artifacts (12.67% Strongly Agree, 55.33% Agree) are the top three cause categories. Some practitioners also suggested 5 types of causes (with 11 instances of causes) that were not part of the initial taxonomy of causes in microservices systems. We added microservices practitioners’ suggested types of causes in Table 5 and the Cause Taxonomy sheet in [19]. One representative quotation about suggested types of causes is depicted below.
@@
Practitioners’ Perspective on Solutions: We also recorded the practitioners’ perspective about the solutions for the issues occurring during microservices system development. The practitioners’ responses presented in Table 5 show that SC1: Add Artifacts (42.00% Strongly Agree, 39.33% Agree), SC3: Modify Artifact (27.33% Strongly Agree, 37.33% Agree), and SC6: Manage Configuration and Execution (18.00% Strongly Agree, 48.67% Agree) and are the top three solution categories that have been used to address the microservices issues. One practitioner also suggested one type of solution (with one instance of solution) that was not part of the initial taxonomy of solutions in microservices systems. We added microservices practitioners’ suggested types of solutions in Table 5 and the Solution Taxonomy sheet in [19]. One representative quotation about the suggested solutions is depicted below.
@@
Table 5: Practitioners’ perspective (in %) on the issue, cause, and solution categories in microservices systems
@@
Statistical significance on the issue, cause, and solution categories in microservices systems: We analyzed the practitioners’ responses across one pair of demographic groups in Table 6. The first column of Table 6 lists the categories of issues, causes, and solutions presented to the survey participants. The subsequent columns of Table 6 show the Likert Distribution, Mean, p-value, and Effect Size. Analyzing survey data using the Likert Distribution, Mean, p-value, and Effect Size provides insights into response patterns, central tendency, statistical significance, and practical significance, respectively. The practitioners’ responses are grouped into Experience $\leq$ 6 Years vs. Experience > 6 Years group to check the test whether there are statistical differences between the two groups on the same variable or not.
@@
Key Findings of RQ4: (i) There are no major statistically significant differences in the issue, cause, and solution categories in practitioners’ responses, (ii) practitioners frequently face Technical Debt, CI/CD, and Security issues, (iii) most causes are associated with General Programming Errors, Fragile Code, and Missing Features and Artifacts, and (iv) most solutions are associated with Add Artifacts, Modify Artifact, and Manage Configuration and Execution. The survey participants generally confirmed the categories of issues, causes, and solutions derived from mining developer discussions in open-source microservices systems. However, the survey participants also indicated several other types of issues, causes, and solutions that help improve our taxonomies.
Table 6: Statistical significance on the issue, cause, and solution categories in microservices systems
@@ 5.1 Internal Validity
Improper project selection. The first internal validity threat to our study is an improper selection of OSS projects for executing our research plan. Mitigation. We used a multi-step project selection approach to control the possible threat associated with subject system selection (see Phase 1 in Section 2.2 and Figure 2). A step-wise and criteria-driven approach for the project selection has been used that helped us to include the relevant (see Table 1) and eliminate irrelevant open-source microservices projects.
@@
Participants selection and background: The survey participants may not have sufficient expertise to answer questions. Mitigation. We searched for microservices practitioners through personal contacts and relevant platforms (e.g., LinkedIn and GitHub). We explicitly made participants’ characteristics (e.g., roles and responsibilities) in the interview and survey preamble, and we selected only those practitioners with sufficient experience designing, developing, and operationalizing microservices systems. For example, the average experience of the interview participants is 5.33 years, and they mainly have responsibilities for designing and developing microservices systems (see Table 2 and Figure 3).
@@ 5.2 Construct Validity
Inadequate explanation of the constructs: This threat refers to the fact that the study constructs are not sufficiently described. Mitigation. To deal with this threat, we prepared the protocols for mining developer discussions and conducting interviews and surveys, and these protocols were continuously improved during the internal meetings, feedback, and taxonomy refinements. Mainly, the authors had meetings (i) to establish a common understanding of issues, causes, and solutions (see Figure 1), (ii) for extracting the required data to answer the RQs, and (iii) for evaluating interview and survey question format, understandability, and consistency. We also invited two survey-based research experts to check the validity and integrity of the survey questions. Based on their feedback, we included Figure 4, Table 3, and Table 4 as a part of the interview and survey questions for improving the participants’ understandability of the issues, their causes and solutions in microservices systems.
@@ 5.3 External Validity
textitInterview data saturation: This threat arises from the possibility that the 15 interviews we conducted to validate and expand our proposed taxonomies may not be sufficient to achieve data saturation in our study. Mitigation. To address this threat, we conducted interviews with practitioners who had varied backgrounds (i.e., responsibilities, skills in programming languages, work domains, experience, countries) in microservices systems (see Table 2). Second, we asked open-ended questions, which allowed the interviewees to spontaneously express their views on the issues, causes, and solutions (see the Interview Questionnaire sheet in [19]). Third, we presented the interviewees with pre-existing taxonomies that were derived from analyzing developer discussions in 15 open-source microservices systems. We asked the interviewees to pinpoint any issue, cause, and solution they felt absent. Fourth, while we were analyzing interview data, we started observing recurring themes and patterns in the responses, which was an indication of data saturation. As we did not encounter any new significant issues, causes, or solutions after the 15th interview, we decided to stop conducting further interviews. However, we encourage further research involving more interviews or additional data sources to provide more insights and refinement to our findings.
@@ 6.1.1 Mining (Anti-)Patterns and Tactics to Architect MSA
From an MSA perspective, architectural patterns [34] and tactics [66] represent empirically derived knowledge, best practices, and recurring solutions to address frequently occurring issues during architecture-centric design and development of service-driven systems [10]. To discover patterns and tactics for architecting MSA, a number of studies have focused on analyzing open source repositories – mining version controls, searching change logs, and exploring design documents etc. – to investigate historical data that represents recurring solutions as patterns [10, 67, 68, 69].
The investigation of historical data involves postmortem analysis of development activities (e.g., software refactoring, testing, evolution) [70] as well as the knowledge of architects (e.g., developer discussions, code documentation) [71], which is captured via open source repositories, such as GitHub, Stack Overflow, or customized databases [10, 67, 71]. Specifically, Marquez et al. [10] explored source code artifacts of microservices projects, i.e., configuration files, and framework dependencies, available on GitHub to mine 17 architectural patterns addressing a set of quality attributes. Compared to the design and implementation-specific knowledge for microservices reported in [10, 67], Armin et al. [68] explored the evolution phase from traditional architectures to MSA and identified 15 migration patterns that support the evolution of legacy software towards a modular MSA. Compared to the pattern-based solutions discussed above, architectural tactics represent design decisions that focus on improving one specific quality aspect of MSA, such as service availability and fault avoidance for security critical systems [66]. Amoroso et al. [72] curated and manually labeled a dataset of 378 Dockerized open-source microservice projects from an initial pool of 389,559 candidates. The dataset includes project metadata such as contributors, dependencies, and application purposes, supporting future empirical research on microservice practices and evolution. Lercher et al. [73] investigated API evolution in microservice systems through interviews with practitioners, identifying six strategies and six challenges in handling REST API changes. They proposed automated mechanisms for extracting structural and behavioral API changes and notifying consumer teams. Their findings highlight the socio-technical complexities of API evolution in microservices, complementing prior work on migration and evolvability. Lubas et al. [74] mined GitHub to curate two datasets: one of 553 microservice applications and another of 8 representative workload repositories. Their analysis provides insights into programming languages, frameworks, and architectural components used in practice, supporting benchmarking and empirical studies on workload characterization in microservice systems.
@@
While patterns promote best practices to develop MSA, anti-patterns represent a class of patterns that have been perceived as best practices and commonly used but are proven to be ineffective and/or counterproductive, such as bad smells in service design [14]. Taibi et al. [70] conducted interviews of microservices developers, i.e., accumulating practitioners’ experiences to create a taxonomy of 20 anti-patterns including organizational (team oriented and technology/tool oriented) and technical (internal and communication) anti-patterns. The anti-pattern taxonomy aims to help microservices practitioners to avoid counterproductive patterns and design decisions. Tighilt et al. [75] proposed a specification and detection approach for microservice antipatterns. Their study defined formal detection rules, validated them on open-source systems, and demonstrated their utility for supporting maintenance and quality improvement in microservice architectures. This work goes beyond descriptive taxonomies by providing automated mechanisms for identifying recurring bad practices. In comparison to the repository mining approaches that investigate code and architectural-centric artifacts available on GitHub such as [66, 10, 69], Bandeira et al. [71] reflected a human-centric view of MSA design by analyzing developer discussions on Stack Overflow. Their study classified a total of 1,043 microservice tagged posts into three categories namely technical (44.87%), conceptual (30%), and non-related (25.13%) discussions to explore the processes, tools, issues, solutions, etc. that developers find most exciting or challenging while implementing MSA.
@@ 6.3 Conclusive Summary
We provide a comparative analysis of the most relevant existing research and our proposed study in Table 7 based on (i) focus of investigation: specific aspects to be empirically investigated, (ii) research method: specific approaches employed for investigation, (iii) number of issues: specific number of the identified issues, and (iv) method for evaluation: specific method for evaluation. Additionally, the study reference, year of publication, and contribution of the research study are provided as complementary information for a fine-grained presentation of the results. For example, an empirical study reported in [15] published in 2021 relies on an industrial survey (i.e., applied research method) to investigate common faults and debugging practices in microservices systems (i.e., focus of investigation). The study results (i.e., main contribution) highlight typical faults, state-of-practice on debugging them, and indicate the challenges faced by developers for microservices systems.
Table 7: Comparative analysis and summary of existing research vs. this research
@@
[15] 2021 Fault analysis, Debugging Industry survey Not explicit mentioned
@@
[16] 2020 Performance issues MSA application monitoring Not explicit mentioned
@@
[76] 2019 Evolvability analysis Structured interviews 16 types of challenges
@@
[79] 2020 Code smells Static code analysis 11 types code smells
@@
[82] 2023 Security points Repository mining
@@
[83] 2022 Microservice problems and solutions Repository mining 14 problems, 47 solutions
@@
[84] 2022 Architectural smells Interviews, Repository mining 6 types of architectural smells
@@
[75] 2023 Antipattern specification and detection Repository mining, Static analysis Catalog of antipatterns
Contribution: This study formalizes the specification of microservice antipatterns and provides automated detection rules validated on open-source systems to support maintenance and quality assurance.
[72] 2024 Dataset of OSS microservice projects Repository mining, Manual curation 378 projects
Contribution: A curated dataset of 378 Dockerized open-source microservice projects.
[73] 2024 API evolution strategies, challenges, and automation Practitioner interviews (17), Proposed automation 6 strategies, 6 challenges
Contribution: An empirical study on strategies and pitfalls in microservice API evolution, with proposals for automated change extraction and notification.
[74] 2025 Dataset of microservice applications and workloads Repository mining, Empirical analysis 553 applications, 8 workloads
Contribution: Curated two GitHub-based datasets of 553 microservice applications and 8 workload repositories, offering insights into languages, frameworks, and components for benchmarking and workload studies.
Our previous study [18] 2022 Issues and Causes in microservices systems Repository mining 138 types of issues, 109 types of causes
@@
This study 2023 Issues, Causes, and Solutions in microservices systems Repository mining (2,641 issues), Structured interviews (15 interviews), Practitioners’ survey (150 responses) 402 types of issues, 228 types of causes, 177 types of solutions
@@ Acknowledgments
This work is partially sponsored by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311. The authors also thank the participants of the interviews and online survey.
Declaration of AI Assistance
During the preparation of this work, the author(s) used ChatGPT to refine grammar, improve sentence structure, and resolve formatting issues. After utilizing this tool, the author(s) thoroughly reviewed and edited the content as needed, taking full responsibility for the final publication.
[…]